# KinetiX: A performance portable code generator for chemical kinetics and transport properties

### v1

@@ Abstract
We present KinetiX, a software toolkit to generate computationally efficient fuel-specific routines for the chemical source term, thermodynamic and mixture-averaged transport properties for use in combustion simulation codes. The C++ routines are designed for high-performance execution on both CPU and GPU architectures. On CPUs, chemical kinetics computations are optimized by eliminating redundant operations and using data alignment and loops with trivial access patterns that enable auto-vectorization, reducing the latency of complex mathematical operations. On GPUs, performance is improved by loop unrolling, reducing the number of costly exponential evaluations and limiting the number of live variables for better register usage. The accuracy of the generated routines is checked against reference values computed using Cantera and the maximum relative errors are below $10^{-5}\%$. We evaluate the performance of the kernels on some of the latest CPU and GPU architectures from AMD and NVIDIA, i.e., AMD EPYC 9653, AMD MI250X, and NVIDIA H100. The routines generated by KinetiX outperform the general-purpose Cantera library, achieving speedups of up to 2.4x for species production rates and 3.2x for mixture-averaged transport properties on CPUs. Compared to the routines generated by PelePhysics (CEPTR), KinetiX achieves speedups of up to 2.6x on CPUs and 1.7x on GPUs for the species production rates kernel on a single-threaded basis.
@@ 2.2.1 Forward and reverse rate constants
The computation of the reverse rate constants $k_{r}$ requires the evaluation of the equilibrium constants $K_{p}$ (Eq. (18)), which involves computing the exponential of the sum of the individual species Gibbs energy $G^{\circ}_{k}$ for each reaction. For the C2H5OH mechanism, 1180 exponentials would have to be evaluated (the reverse rate constants of irreversible reactions are identically zero). If the exponential of the sum is written as a product of the exponentials of each species, the individual terms can be evaluated using Eq. (19), where all $a_{i,k},i=0,\cdots 6$ are known parameters from the reaction mechanism and can be stored and aligned in densely packed arrays to enable auto-vectorization. Once $G^{\circ}_{k}$, the Gibbs energy of the species, has been computed, the exponentials can be evaluated in a vectorized manner (see Fig. 2(a) lines 40-41). Using this approach, only 129 exponentials are evaluated for $k_{r}$ in the ethanol mechanism, compared to 1180 exponentials that would be computed in the generated code presented by Zirwes et al. [11]. In addition, by pre-computing the reciprocal Gibbs exponentials for the reactants, costly division operations can be replaced with more efficient multiplications in the final calculation for each reaction. Tests on a single AMD EPYC 7763 CPU showed that reducing the number of exponential evaluations for computing reverse rate constants can lead to a 1.5x speedup in the species production kernel for the ethanol mechanism. However, the magnitude of the speed-up depends on the mechanism complexity. For instance, with a smaller mechanism like GRI-Mech 3.0 [31] with 53 species and 325 reactions the speed-up is approximately 1.25x.
@@ 2.3.2 Mixture-averaged diffusion coefficients
One way to speed-up the computation of the mixture-averaged diffusion coefficients on GPUs is to evaluate the reciprocal polynomial directly (Fig. 3(b)) and leverage the symmetry of $D_{kj}$. While this can result in significant speed up by minimizing redundant calculations, it also increases the register pressure and can lead to register spilling, low occupancy, and underutilization of mathematical units. Similar to the CPU implementation, another approach can be used that does not assume symmetry and evaluates each entry of the matrix completely. This effectively reduces the number of intermediate values at the cost of doubling the number of polynomial evaluations. The complete evaluation is particularly suitable for large mechanisms since the number of intermediate values that need to be stored in the symmetric evaluation scales quadratically with the number of species.
Figure 3: Mixture-average diffusivities code generation
@@ 3.1 Validation
The accuracy of the generated kernels is automatically checked against reference data, which are precomputed using Cantera using the Jupyter notebook that simulates autoignition in a constant pressure reactor. Three thermochemical states of pre-ignition, ignition and post-ignition are chosen, and the notebook computes and stores the thermodynamic data together with the reaction rates, the net production rates, and mixture-averaged transport properties.
Table 3 reports the values for the ignition case. The thermodynamic properties and reaction rates show negligible mean and maximum relative differences. In the mixture-averaged transport properties, the relative differences are moderately higher. The discrepancy was found to be caused by the fitting algorithms used to fit the expressions for the pure species properties. In KinetiX we use the polyfit function of NumPy, while Cantera employs the Eigen library [39]. The algorithmic differences between the two libraries lead to small differences in the fitted coefficients, which lead to slightly larger errors in the evaluation of the mixture-averaged properties. It is worth mentioning that the relative error between the transport properties of the pure species computed internally in KinetiX and the reference values from Cantera is in general smaller than $10^{-12}\%$, and the larger errors in the mixture-averaged values are primarily due to the different polynomial fitting functions. Nevertheless, the maximum relative errors for the mixture-averaged transport properties are below $10^{-5}\%$.
Table 3: Summary of difference between KinetiX-generated properties and Cantera reference values. Error statistics are based on the mean and maximum relative error percentages $E_{rel}$ for each property.
Model Properties Mean (%) Maximum (%)
H2 Thermodynamic properties 1.987 × 10-14 4.220 × 10-14
Reaction rates 1.447 × 10-9 8.677 × 10-9
Transport properties 3.765 × 10-8 1.800 × 10-6
GRI-Mech 3.0 Thermodynamic properties 2.623 × 10-14 7.681 × 10-14
Reaction rates 6.160 × 10-11 1.899 × 10-9
Transport properties 1.365 × 10-7 1.126 × 10-6
C2H5OH Thermodynamic properties 7.292 × 10-15 6.015 × 10-14
Reaction rates 3.629 × 10-9 2.329 × 10-7
Transport properties 2.826 × 10-8 9.874 × 10-6
@@ 3.2 Performance analysis
For comparison, the performance of the kernels using CEPTR-generated routines was assessed. However, it is important to note that CEPTR only generates the fitting coefficients for the transport quantities, while the mixture-averaged properties are evaluated internally. Furthermore, CEPTR employs a third-degree polynomial for the fit and evaluates the logarithm of the transport properties of each species, adhering to the approach of CHEMKIN. In KinetiX, we use a fourth-degree polynomial and evaluate the species transport properties directly, following the approach of Cantera. These differences precluded a direct comparison of transport kernel performance between KinetiX and CEPTR. The CPU routines are also compared with Cantera using the same metrics.
@@ 3.2.1 Species production rates
In contrast, CPU performance exhibits higher initial throughput values and reaches its peak performance at almost two orders of magnitude smaller number of thermochemical states than the GPUs. The earlier saturation point underscores the CPU ability to achieve optimal performance on smaller problem sizes. These observations highlight a crucial consideration when evaluating the performance of such kernels across different computing platforms, namely that there is an optimal performance window with respect to problem size (blue and red shaded areas in Fig. 5) for which peak throughput is achieved. While GPUs generally offer superior performance for large-scale problems, CPUs can actually achieve higher throughput for smaller problem sizes.
@@
For the GRI 3.0 mechanism, there is a significant decrease in throughput on both GPUs, which is due to the significantly higher register pressure resulting from the 178 live double-precision variables, i.e. 1376 bytes of information that needs to be stored. Even with the maximum number of registers available per thread (255), 404 bytes of data is spilled to off-chip global memory. In addition, this high register pressure leads to lower occupancy and under-utilization of mathematical units. The NVIDIA H100 GPU features a large L2 cache of 50 MB. While spilled data is initially stored in global memory, the frequent accesses to this data can lead to it being cached in the L2, which can significantly reduce latency compared to repeatedly fetching the data directly from global memory. The large cache size increases the likelihood that frequently accessed spilled data remains in cache, potentially offering substantial performance benefits. In contrast, the AMD MI250X, with an L2 cache size of 16 MB, presents more of a challenge. The limited cache memory exacerbates the effects of the data spill, significantly decreasing throughput. The difference in performance between the two GPUs is also significantly smaller for the CH4 mechanism.
@@
When comparing the performance between KinetiX and CEPTR on GPUs (red and green bars, respectively, in Fig. 5), we observe that KinetiX consistently performs better for all mechanisms. Both codes utilize an approach where reactions are unrolled and computed sequentially on GPUs with similar number of live variables during kernel execution. The primary distinction between the two codes lies in the reduction of exponential evaluations described in Sec. 2. By pre-computing the Gibbs exponentials, KinetiX significantly reduces the number of evaluations required to compute the reverse rate constants, which scales with the ratio of the number of reactions to the number of species. This reduction becomes increasingly beneficial for larger mechanisms. Similarly, the reduction of exponential evaluations in Arrhenius law (Sec. 2) is proportional to the mechanism size, making the KinetiX optimization strategy more effective with increasing mechanism size. For hydrogen, the performance difference is small, with both codes achieving similar throughput. The H2 mechanism, with 21 reactions and 9 species, allows for only 12 avoided exponential evaluations at the cost of 21 additional divisions, which negatively impact performance. Moreover, only two reactions benefit from reduced exponentials in Arrhenius law. Consequently, the overall performance difference between KinetiX and CEPTR remains small on both GPUs. For methane, the performance difference increases, with KinetiX achieving a speedup of up to 1.49x on the H100 GPU and 1.16x on the MI250X GPU. This improvement is due to the mechanism having a higher reaction-to-species ratio and a larger number of exponentials falling into the strategies described in Sec. 2. The ethanol mechanism reinforces this trend, with the performance benefit of KinetiX increasing further to 1.69x on the H100 GPU and 1.27x on the MI250X GPU.
@@ 3.2.2 Mixture-averaged transport properties
Figure 7 shows the maximum achieved throughput in GDOF/s for KinetiX on GPUs and CPUs and Cantera on CPUs. The transport kernels on GPUs exhibit behavior similar to that of the species production rates kernels. For hydrogen, the throughput is high, with AMD MI250X achieving higher values compared to H100. In the transport kernel, the number of live variables can scale quadratically with the number of species, often exceeding the small on-chip memory allocated to each thread, even for small mechanisms, resulting in register spilling, and low occupancy. However, by using techniques such as loop unrolling, computing the contributions of individual species to the mixture-averaged viscosity, and evaluating the complete diffusivities matrix (Sec. 2.3.2), the number of live variables in the transport kernels can be significantly reduced. As a result, these optimized transport kernels require fewer live variables compared to the species production rates kernel. For hydrogen, the memory requirements per thread can easily be accommodated within the limit of 255 registers per thread of each GPU, at a value of 50 registers per thread. The lower number of registers results in higher occupancy compared to the production rates kernel. The larger register file size of the AMD MI250X allows more threads per CU to remain active, which in combination with the higher number of CUs, leads to greater theoretical parallelism for the H2 mechanism and higher throughput. For the GRI-Mech 3.0 mechanism, the larger number of live variables results in higher register pressure, lower occupancy and data spillage to off-chip memory. The H100 increased L2 cache, which is more than three times the size of the AMD MI250X, allows it to manage this spilled data more effectively, resulting in better performance. For the C2H5OH mechanism, the transport kernel becomes global memory-bound on both GPU architectures, resulting in significantly lower throughput.
@@
It should be pointed out that unlike the GRXN/s metric used for species production rates, the GDOF/s metric used for mixture-averaged transport properties is not independent of the problem size. This is because the number of floating point operations for transport properties scales quadratically with the number of species. Consequently, the throughput values decrease with increasing mechanism size, even on CPUs. This effect is observed despite the large cache size of the AMD EPYC 9654 CPU, which can accommodate even the largest mechanism. In comparison to Cantera, the throughput improvement becomes more pronounced with larger mechanism sizes. For mixture-averaged transport properties, the coefficients for the polynomial fits of the pure species properties are densely packed in memory and defined as compile-time constants, which together with loop optimization strategies enables the auto-vectorization of the loops for a more efficient computation of the mixture-averaged transport properties. For the largest mechanism, a significant speed-up of 3.2x is achieved compared to Cantera. This demonstrates the effectiveness of our optimization techniques on CPUs, particularly for complex combustion kinetics.
@@ 4 Conclusions
In order to run KinetiX on different computing architectures, the OCCA library, which allows for runtime code generation for different threading programming paradigms was used to create kernels that call the generated C++ routines. The kernel performance was evaluated on some of the latest CPU and GPU architectures from AMD and NVIDIA i.e., AMD EPYC 9653, AMD MI250X and NVIDIA H100. Performance benchmarking showed that the routines generated by KinetiX outperform the general-purpose Cantera library with speed-ups of up to 2.4x for species production rates and 3.2x for mixture-averaged transport properties on CPU. Compared to CEPTR, KinetiX achieves speed-ups of up to 2.6x on CPU and up to 1.7x on GPUs for the species production rates kernel on a single-threaded basis. The most significant performance improvements were observed with the largest reaction mechanism (C2H5OH), whereas the smallest mechanism tested (H2) yielded more modest gains.
Planned development work for KinetiX includes support for pressure-dependent reaction formulations based on logarithmic or Chebyshev polynomial interpolation and quasi-steady-state approximation (QSSA) species. Further studies will focus on exploring different parallelization strategies for GPUs. Currently, the per-thread or grid-based parallelization approach used in KinetiX is effective for small to medium sized kinetic problems. However, as HPC trends move towards GPUs with larger register count and cache size, these limitations may be mitigated in the future, enabling the efficient computation of larger mechanisms. Furthermore, the reaction mechanisms for carbon-free fuels generally involve fewer species and reactions. Nonetheless, exploring alternative strategies, such as collaborative thread execution or warp specialization techniques, would be valuable to determine the best performance scaling with reaction mechanism size on GPUs. In addition, we will explore mixed precision formulations for the generated routines which can potentially offer significant performance on GPUs.

### v2

@@ Abstract
We present KinetiX, a software toolkit to generate computationally efficient fuel-specific routines for the chemical source term, thermodynamic and mixture-averaged transport properties for use in combustion simulation codes. The C++ routines are designed for high-performance execution on both CPU and GPU architectures. On CPUs, chemical kinetics computations are optimized by eliminating redundant operations and using data alignment and loops with trivial access patterns that enable auto-vectorization, reducing the latency of complex mathematical operations. On GPUs, performance is improved by loop unrolling, reducing the number of costly exponential evaluations and limiting the number of live variables for better register usage. The accuracy of the generated routines is checked against reference values computed using Cantera and the maximum relative errors are below $10^{-7}$. We evaluate the performance of the kernels on some of the latest CPU and GPU architectures from AMD and NVIDIA, i.e., AMD EPYC 9653, AMD MI250X, and NVIDIA H100. The routines generated by KinetiX outperform the general-purpose Cantera library, achieving speedups of up to 2.4x for species production rates and 3.2x for mixture-averaged transport properties on CPUs. Compared to the routines generated by PelePhysics (CEPTR), KinetiX achieves speedups of up to 2.6x on CPUs and 1.7x on GPUs for the species production rates kernel on a single-threaded basis.
PROGRAM SUMMARY
Program Title: KinetiX
Developer’s repository link: https://github.com/bogdandanciu/KinetiX
Licensing provisions: BSD 2-clause
Programming language: Python, C++
Nature of problem: Combustion simulations require efficient computation of chemical source terms, thermodynamic and transport properties for diverse fuel types. The challenge is optimizing these computations for both CPUs and GPUs without compromising accuracy.
Solution method: Starting from an input file containing kinetic parameters, thermodynamic and transport data, KinetiX generates fuel-specific routines to compute species production rates, thermodynamic and mixture-averaged transport properties for high-performance execution on both CPU and GPU architectures.
@@ 2.2.1 Forward and reverse rate constants
The computation of the reverse rate constants $k_{r}$ requires the evaluation of the equilibrium constants $K_{p}$ (Eq. (18)), which involves computing the exponential of the sum of the individual species Gibbs energy $G^{\circ}_{k}$ for each reaction. For the C2H5OH mechanism, 1180 exponentials would have to be evaluated (the reverse rate constants of irreversible reactions are identically zero). If the exponential of the sum is written as a product of the exponentials of each species, the individual terms can be evaluated using Eq. (19), where all $a_{i,k},i=0,\cdots 6$ are known parameters from the reaction mechanism and can be stored and aligned in densely packed arrays to enable auto-vectorization. Once $G^{\circ}_{k}$, the Gibbs energy of the species, has been computed, the exponentials can be evaluated in a vectorized manner (see Fig. 2(a) lines 40-41). Using this approach, only 129 exponentials are evaluated for $k_{r}$ in the ethanol mechanism, compared to 1180 exponentials that would be computed in the generated code presented by Zirwes et al. [11]. In addition, by pre-computing the reciprocal Gibbs exponentials for the reactants, costly division operations can be replaced with more efficient multiplications in the final calculation for each reaction. Tests on a single AMD EPYC 7763 CPU showed that reducing the number of exponential evaluations for computing reverse rate constants can lead to a 1.5x speedup in the species production kernel for the ethanol mechanism. However, the magnitude of the speedup depends on the mechanism complexity. For instance, with a smaller mechanism like GRI-Mech 3.0 [31] with 53 species and 325 reactions the speedup is approximately 1.25x.
@@ 2.3.2 Mixture-averaged diffusion coefficients
One way to speedup the computation of the mixture-averaged diffusion coefficients on GPUs is to evaluate the reciprocal polynomial directly (Fig. 3(b)) and leverage the symmetry of $D_{kj}$. While this can result in significant speed up by minimizing redundant calculations, it also increases the register pressure and can lead to register spilling, low occupancy, and underutilization of mathematical units. Similar to the CPU implementation, another approach can be used that does not assume symmetry and evaluates each entry of the matrix completely. This effectively reduces the number of intermediate values at the cost of doubling the number of polynomial evaluations. The complete evaluation is particularly suitable for large mechanisms since the number of intermediate values that need to be stored in the symmetric evaluation scales quadratically with the number of species.
Figure 3: Mixture-average diffusivities code generation.
@@ 3.1 Validation
The accuracy of the generated kernels is automatically checked against reference data, which are precomputed using Cantera and a Jupyter notebook that simulates autoignition in a constant pressure reactor. Three thermochemical states of pre-ignition, ignition and post-ignition are chosen, and the notebook computes and stores the thermodynamic data together with the reaction rates, the net production rates, and mixture-averaged transport properties.
Table 3 reports the values for the ignition case. The thermodynamic properties and reaction rates show negligible mean and maximum relative differences. In the mixture-averaged transport properties, the relative differences are moderately higher. The discrepancy was found to be caused by the fitting algorithms used to fit the expressions for the pure species properties. In KinetiX we use the polyfit function of NumPy, while Cantera employs the Eigen library [39]. The algorithmic differences between the two libraries lead to small differences in the fitted coefficients, which lead to slightly larger errors in the evaluation of the mixture-averaged properties. It is worth mentioning that the relative error between the transport properties of the pure species computed internally in KinetiX and the reference values from Cantera is in general smaller than $10^{-14}$, and the larger errors in the mixture-averaged values are primarily due to the different polynomial fitting functions. Nevertheless, the maximum relative errors for the mixture-averaged transport properties are below $10^{-7}$.
Table 3: Summary of difference between KinetiX-generated properties and Cantera reference values. Error statistics are based on the mean and maximum relative error $E_{rel}$ for each property.
Model Properties Mean Maximum
H2 Thermodynamic properties 1.987 × 10-16 4.220 × 10-16
Reaction rates 1.447 × 10-11 8.677 × 10-11
Transport properties 3.765 × 10-10 1.800 × 10-8
GRI-Mech 3.0 Thermodynamic properties 2.623 × 10-16 7.681 × 10-16
Reaction rates 6.160 × 10-13 1.899 × 10-11
Transport properties 1.365 × 10-9 1.126 × 10-8
C2H5OH Thermodynamic properties 7.292 × 10-17 6.015 × 10-16
Reaction rates 3.629 × 10-11 2.329 × 10-9
Transport properties 2.826 × 10-10 9.874 × 10-8
@@ 3.2 Performance analysis
For comparison, we evaluated the performance against both CEPTR and Cantera using the same metrics. While a direct comparison of species production rates was possible since all three packages implement the same formulations for chemical kinetics (detailed in A), the comparison of transport properties required special consideration. CEPTR follows the CHEMKIN approach by employing third-degree polynomials and evaluating the logarithm of the transport properties, whereas both KinetiX and Cantera use fourth-degree polynomials and evaluate the properties directly. In addition, CEPTR only generates the fitting coefficients for the transport quantities, with mixture-averaged properties evaluated internally. These differences in transport property computation precluded a direct performance comparison between KinetiX and CEPTR for the transport kernel, although we were able to compare CPU performance with Cantera due to their similar approaches.
@@ 3.2.1 Species production rates
In contrast, CPU performance exhibits higher initial throughput values and reaches its peak performance at almost two orders of magnitude smaller number of thermochemical states than the GPUs. The earlier saturation point underscores the CPU’s ability to achieve optimal performance on smaller problem sizes. These observations highlight a crucial consideration when evaluating the performance of such kernels across different computing platforms, namely that there is an optimal performance window with respect to problem size (blue and red shaded areas in Fig. 4) for which peak throughput is achieved. While GPUs generally offer superior performance for large-scale problems, CPUs can actually achieve higher throughput for smaller problem sizes.
@@
For the GRI 3.0 mechanism, there is a significant decrease in throughput on both GPUs, which is due to the significantly higher register pressure resulting from the 178 live double-precision variables, i.e. 1376 bytes of information that needs to be stored. Even with the maximum number of registers available per thread (255), 404 bytes of data are spilled to off-chip global memory. In addition, this high register pressure leads to lower occupancy and under-utilization of mathematical units. The NVIDIA H100 GPU features a large L2 cache of 50 MB. While spilled data is initially stored in global memory, the frequent accesses to this data can lead to it being cached in the L2, which can significantly reduce latency compared to repeatedly fetching the data directly from global memory. The large cache size increases the likelihood that frequently accessed spilled data remains in cache, potentially offering substantial performance benefits. In contrast, the AMD MI250X, with an L2 cache size of 16 MB, presents more of a challenge. The limited cache memory exacerbates the effects of the data spill, significantly decreasing throughput. The difference in performance between the two GPUs is also significantly smaller for the CH4 mechanism.
@@
When comparing the performance between KinetiX and CEPTR on GPUs (red and green bars, respectively, in Fig. 5), we observe that KinetiX consistently performs better for all mechanisms. Both codes utilize an approach where reactions are unrolled and computed sequentially on GPUs with similar number of live variables during kernel execution. The primary distinction between the two codes lies in the reduction of exponential evaluations described in Sec. 2. By pre-computing the Gibbs exponentials, KinetiX significantly reduces the number of evaluations required to compute the reverse rate constants, which scales with the ratio of the number of reactions to the number of species. This reduction becomes increasingly beneficial for larger mechanisms. Similarly, the reduction of exponential evaluations in Arrhenius law (Sec. 2) is proportional to the mechanism size, making the KinetiX optimization strategy more effective with increasing mechanism size. For hydrogen, the performance difference is small, with both codes achieving similar throughput. The H2 mechanism, with 9 species and 21 reactions, allows for only 12 avoided exponential evaluations at the cost of 21 additional divisions, which negatively impact performance. Moreover, only two reactions benefit from reduced exponentials in Arrhenius law. Consequently, the overall performance difference between KinetiX and CEPTR remains small on both GPUs. For methane, the performance difference increases, with KinetiX achieving a speedup of up to 1.49x on the H100 GPU and 1.16x on the MI250X GPU. This improvement is due to the mechanism having a higher reaction-to-species ratio and a larger number of exponentials falling into the strategies described in Sec. 2. The ethanol mechanism reinforces this trend, with the performance benefit of KinetiX increasing further to 1.69x on the H100 GPU and 1.27x on the MI250X GPU.
@@ 3.2.2 Mixture-averaged transport properties
Figure 7 shows the maximum achieved throughput in GDOF/s for KinetiX on GPUs and CPUs and Cantera on CPUs. The transport kernels on GPUs exhibit behavior similar to that of the species production rates kernels. For hydrogen, the throughput is high, with MI250X achieving higher values compared to H100. In the transport kernel, the number of live variables can scale quadratically with the number of species, often exceeding the small on-chip memory allocated to each thread, even for small mechanisms, resulting in register spilling, and low occupancy. However, by using techniques such as loop unrolling, computing the contributions of individual species to the mixture-averaged viscosity, and evaluating the complete diffusivities matrix (Sec. 2.3.2), the number of live variables in the transport kernels can be significantly reduced. As a result, these optimized transport kernels require fewer live variables compared to the species production rates kernel. For hydrogen, the memory requirements per thread can easily be accommodated within the limit of 255 registers per thread of each GPU, at a value of 50 registers per thread. The lower number of registers results in higher occupancy compared to the production rates kernel. The larger register file size of the MI250X allows more threads per CU to remain active, which in combination with the higher number of CUs, leads to greater theoretical parallelism for the H2 mechanism and higher throughput. For the GRI-Mech 3.0 mechanism, the larger number of live variables results in higher register pressure, lower occupancy and data spillage to off-chip memory. The H100’s increased L2 cache size, which is more than three times the size of the MI250X, allows it to manage this spilled data more effectively, resulting in better performance. For the C2H5OH mechanism, the transport kernel becomes global memory-bound on both GPU architectures, resulting in significantly lower throughput.
@@
It should be pointed out that unlike the GRXN/s metric used for species production rates, the GDOF/s metric used for mixture-averaged transport properties is not independent of the problem size. This is because the number of floating point operations for transport properties scales quadratically with the number of species. Consequently, the throughput values decrease with increasing mechanism size, even on CPUs. This effect is observed despite the large cache size of the AMD EPYC 9654 CPU, which can accommodate even the largest mechanism. In comparison to Cantera, the throughput improvement becomes more pronounced with larger mechanism sizes. For mixture-averaged transport properties, the coefficients for the polynomial fits of the pure species properties are densely packed in memory and defined as compile-time constants, which together with loop optimization strategies enables the auto-vectorization of the loops for a more efficient computation of the mixture-averaged transport properties. For the largest mechanism, a significant speedup of 3.2x is achieved compared to Cantera. This demonstrates the effectiveness of our optimization techniques on CPUs, particularly for complex combustion kinetics.
@@ 4 Conclusions
In order to run KinetiX on different computing architectures, the OCCA library, which allows for runtime code generation for different threading programming paradigms was used to create kernels that call the generated C++ routines. The kernel performance was evaluated on some of the latest CPU and GPU architectures from AMD and NVIDIA i.e., AMD EPYC 9653, AMD MI250X and NVIDIA H100. Performance benchmarking showed that the routines generated by KinetiX outperform the general-purpose Cantera library with speedups of up to 2.4x for species production rates and 3.2x for mixture-averaged transport properties on CPU. Compared to CEPTR, KinetiX achieves speedups of up to 2.6x on CPU and up to 1.7x on GPUs for the species production rates kernel on a single-threaded basis. The most significant performance improvements were observed with the largest reaction mechanism (C2H5OH), whereas the smallest mechanism tested (H2) yielded more modest gains.
Planned development work for KinetiX includes support for quasi-steady-state approximation (QSSA) species and analytical Jacobian computation. Further studies will focus on exploring different parallelization strategies for GPUs. Currently, the per-thread or grid-based parallelization approach used in KinetiX is effective for small to medium sized kinetic problems. However, as HPC trends move towards GPUs with larger register count and cache size, these limitations may be mitigated in the future, enabling the efficient computation of larger mechanisms. Furthermore, the reaction mechanisms for carbon-free fuels generally involve fewer species and reactions. Nonetheless, exploring alternative strategies, such as collaborative thread execution or warp specialization techniques, would be valuable to determine the best performance scaling with reaction mechanism size on GPUs. In addition, we will explore mixed precision formulations for the generated routines which can potentially offer significant performance on GPUs.